OpenStudy (anonymous):

HELP PLEASE!!! FAN&&MEDAL!!!! Which pair of angles has congruent values for the sin x° and the cos y°? 70°; 160° 70°; 70 70°; 120° 70°; 20° @EclipsedStar @pooja195 @bohotness @Schen421

OpenStudy (mathmate):

Hint: sin(x)=cos(90-x) for all real values of x.

OpenStudy (bohotness):

looks like math got this :p

OpenStudy (anonymous):

yea ik that but im confused @mathmate @bohotness

OpenStudy (bohotness):

okay

OpenStudy (mathmate):

The pairs of angles are (70°; 160°) ( 70°; 70) ( 70°; 120°) ( 70°; 20°) For example, for the first case, x=70°, y=160°, The questions wants to know if sin(x)=cos(y), or sin(70°)=cos(160°). If they are equal, then answer is choice 1. If not, look at the others.

OpenStudy (anonymous):

so the answer is B? SInce its congruent and congruent means everything is equal. @mathmate

OpenStudy (anonymous):

am I right? @pooja195 @mathmate

OpenStudy (mathmate):

The questions wants to know if sin(x)=cos(y), which means if sin(70)=cos(70). What do you think?

OpenStudy (anonymous):

Do u mind helping me with 3 more please

OpenStudy (mathmate):

If you post new questions, you get more attention, and perhaps more help. I think the rule is one question per post.

OpenStudy (mathmate):

Tag me after your posts if you wish.